John`s income drops from 100.000 € to 90.000 €, so now he takes 3 holiday instead of What is his YED fo holidays?


1
Expert's answer
2022-01-20T12:05:37-0500

Income Elasticity of demand is the the measure of change in quantity demanded due to change in come.

"YED=" "\frac{\Delta D/D}{\Delta I/I}"

D=5

"\Delta D=5-3=2"

"\Delta I=100,000-90,000= 10,000"


="\frac{2/5}{10,000/100,000}"

="\frac{2}{5}\times \frac{100000}{10,000}" = 4